OpenStudy (anonymous):

Julie and Eric row their boat (at a constant speed) 60 miles downstream for 6 hours, helped by the current. Rowing at the same rate, the trip back against the current takes 10 hours. Find the rate of the current.

OpenStudy (anonymous):

down stream she was going \[\frac{60miles}{6hours}=10mph\] while going upstream at :\[\frac{60miles}{10 hours}=6mph\] since we are looking for the current all we have to do is subtract the difference in the speed to get our answer! \[10mph - 6mph = ? mph\]

OpenStudy (anonymous):

4mph

OpenStudy (anonymous):

yep so the current was 4mph!

OpenStudy (anonymous):

it makes sense to me, but my option are 3mph, 1.5, 2, or 8 mph

OpenStudy (anonymous):

yea i got mixed up with this :O its should be like D=RT R=D/T (R+C)=60/6 (R+C)=10 MPH (R-C)=60/10 (R-C)=4 MPH (10-6)/2=4/2=2 MPH IS THE CURRENT

OpenStudy (anonymous):

forgot to divide by 2 :(

OpenStudy (anonymous):

lol ok i was thinking two. makes more sense now! thank you! :)

OpenStudy (anonymous):

no problem! :)